OpenStudy (anonymous):

if x mod y is the remainder when x is divided by y,then (57 mod 6)-(5 mod 4)=?

Parth (parthkohli):

Do you know how to start with this problem?

OpenStudy (anonymous):

no i dont know :(

Parth (parthkohli):

Hmm... do you know what "remainder" means?

OpenStudy (anonymous):

yes i do

Parth (parthkohli):

So can you calculate the remainder when 57 is divided by 6?

OpenStudy (anonymous):

yes the remainder wud be 1

Parth (parthkohli):

No... no.

Parth (parthkohli):

You can calculate the remainder showing long division.

OpenStudy (anonymous):

sorry it wud be 3

Parth (parthkohli):

That's correct, and how about the remainder when you divide 5 by 4?

OpenStudy (anonymous):

then what to do?

Parth (parthkohli):

Do you know what the remainder is when 5 is divided by 4? I'm guiding you to your solution, don't worry. :)

OpenStudy (anonymous):

yes 1.25

OpenStudy (anonymous):

reply plz

OpenStudy (anonymous):

when 5 is divided by 4 remainder is 1

OpenStudy (anonymous):

3-1=2.

OpenStudy (anonymous):

5=4*1+1 dividend=divisor*quotient+remainder

Parth (parthkohli):

What in the world?

Parth (parthkohli):

Yes, it is 3 - 1 = 2.

OpenStudy (anonymous):

yes but the answer wud be 1.25

Parth (parthkohli):

@nitz: No, that doesn't apply everywhere.

OpenStudy (anonymous):

if x mod y is the remainder when x is divided by y,then (57 mod 6)-(5 mod 4)=?

OpenStudy (anonymous):

how is the answer 1.25? You are subtracting the remainders

OpenStudy (anonymous):

where it doesnt apply

OpenStudy (anonymous):

an example

Parth (parthkohli):

@laraib You have to find the remainder when 5 is divided by 4.